# From homeostasis to resource sharing: Biologically and economically aligned multi-objective multi-agent AI safety benchmarks

### v3

@@ 3.1.2 Handwritten Rules Agent (no learning)
This agent type represents the approximate best scores an agent is expected to achieve on the benchmarks. This agent is based on the simulation of hand-written heuristic action rules for a particular benchmark setup. This agent does not have a capability to learn.
@@ 3.1.3 OpenAI Stable Baselines 3 algorithms
We used OpenAI implementations, except for the CNN-based feature detector part, where we had to use our custom CNN class, because the OpenAI implementation required a bigger input image than our environment provided.
@@ 3.1.4 LLM agent
During training, the observations provided to the model included rewards. During the test, the rewards were not revealed to the model. The message history was carried over from train to test.
@@ 3.2 Experimental results
Since the above-mentioned "Random Agent" and "Handwritten Rules Agent" do not learn, the benchmarks did not contain a training phase for these agent types. For these agents, we measured the performance only during the testing phase for 100 episodes, each episode consisting of 400 steps.
OpenAI Stable Baselines 3 algorithms were trained for 2500 episodes and tested for 100 episodes, each episode consisting of 400 steps. Due to compute constraints, there were no trials (model resets).
Due to budget reasons, the LLM agent was trained for only 1 episode and then tested also for 1 episode per each benchmark, again with no trials.
@@
Table 2: Benchmark score per agent type
@@ 3 CNN
0.85 19.88 19.88 19.88 11.74 6.38 14.53 1 1 0.6 19.95 19.9
Danger Tiles -1.15 19.88 19.87 19.49 1.78 5.87 -1.11 -1.79 0 -1.6 19.85 19.89
Predators -3.81 19.2 19.19 19.09 -47.72 -47.72 -29.58 -52.61 -50.24 1.87 -67.25 18.19
Homeostasis -90.75 -92.2 -92.04 -90.77 -92.04 -92.2 -92.55 -92.75 -92.04 -92.75 -92.75 8.2
Sustainability 9.04 1.53 2.79 0.84 0.59 1.59 3.45 0.18 1.01 0.14 1.05 0.64 *
@@
** PPO with weight sharing
*** PPO without weight sharing

### v4

@@ 3.1.2 Handwritten Rules Agent (no learning)
This agent type represents the approximate best scores that an agent is expected to achieve on the benchmarks. This agent is based on the simulation of hand-written heuristic action rules for a particular benchmark setup. This agent does not have a capability to learn.
@@ 3.1.3 OpenAI Stable Baselines 3 algorithms
We used OpenAI implementations, except for the CNN-based feature detector part, where we had to use our custom CNN class because the OpenAI implementation required a bigger input image than our environment provided.
@@ 3.1.4 LLM agent
Training consisted of in-context learning - no fine-tuning was performed. During training, the observations provided to the model included rewards. During the test, the rewards were not revealed to the model. The message history was carried over from train to test.
@@ 3.2 Experimental results
Since the above-mentioned "Random Agent" and "Handwritten Rules Agent" do not learn, the benchmarks did not contain a training phase for these agent types. For these agents, we measured the performance only during the testing phase for 100 trials, each trial consisting of 1 episode, and each episode consisting of 400 steps.
OpenAI Stable Baselines 3 algorithms were tested on 100 trials (model resets). For each trial the model was trained for 1000000 (1M) steps (episode lengths were variable and determined by SB3 algorithm) and after training tested for 1 episode consisting of 400 steps. The model was trained and tested for each benchmark separately.
Due to budget reasons, the LLM agent was tested for only 1 trial per each benchmark. A trial consisted of 1 episode for training and then 1 episode for testing (episodes were 400 steps each).
In the Food Sharing benchmark, the score is averaged over two agents.
@@
Table 2: Average benchmark score per agent type
@@ 3 CNN
0.85 19.88 19.86 19.88 9.46 4.47 6.14 0.50 0.57 0.62 19.95 19.90
Danger Tiles -1.15 19.71 19.81 19.47 10.06 3.99 -0.98 -0.93 -0.83 -1.22 19.85 19.89
Predators -3.81 19.16 19.08 19.00 -39.77 -46.44 -22.12 -38.82 -49.22 -13.89 -67.25 18.19
Homeostasis -90.75 -91.56 -91.22 -90.68 -92.33 -92.35 -92.28 -92.33 -92.30 -92.29 -92.75 8.20
Sustainability 9.04 1.54 1.91 1.03 1.38 1.11 4.67 0.85 0.20 1.66 1.05 0.64 *
@@
** PPO without weight sharing
*** PPO with weight sharing